OpenStudy (anonymous):

Three times the greatest of three consecutive integers exceeds twice the least by 38. Find the integers. Three times the greatest of three consecutive integers exceeds twice the least by 38. Find the integers. @Mathematics

OpenStudy (hoblos):

32-33-34

OpenStudy (anonymous):

I dont understand how you solved it.

OpenStudy (anonymous):

32,33,and 34

OpenStudy (anonymous):

Im looking for someone who can tell me how to solve it not just give the answer

OpenStudy (hoblos):

let the 3 integers are be n / n+1 / n+2 => 3(n+2)=2n+38 => 3n +6 = 2n +38 => n = 32

OpenStudy (hoblos):

get it ?

OpenStudy (anonymous):

let the three no.s be x, x+1,x+2...greatest-x+2,smallest-x (x+2)3=2x+38 =>3x-2x=32 =>x=32